OpenStudy (anonymous):

A potter's wheel, with rotational inertia 46 kg∙m2, is spinning freely at 40 rpm. The potter drops a lump of clay onto the wheel, where it sticks a distance 1.2 m from the rotational axis. If the subsequent angular speed of the wheel and clay is 32 rpm what is the mass of the clay?

OpenStudy (vincent-lyon.fr):

Tip: Use conservation of angular momentum.

OpenStudy (anonymous):

I see the formula is L = r*m*v, but I'm confused about what is L in my question

OpenStudy (anonymous):

Initial angular momentum = final angular momentum u knw L is also = I*W

OpenStudy (anonymous):

46 * 40 = ( 46 + m* (1.2)^2 ) * 32 find m

OpenStudy (anonymous):

Thanks for your help yahoo!